# Co-designing an AI Impact Assessment Report Template with AI Practitioners and AI Compliance Experts

### v1

@@ Discussion
Differences with Existing Templates
We compared our template with the Algorithmic Impact Assessment template from NIST (National Institute of Standards and Technology 2023b), Microsoft’s Responsible AI Impact Assessment Template (Microsoft 2022b), Credo AI’s Standardized Risk Profile (Sherman and Eisenberg 2024), and the Algorithmic Impact Assessment from the Ada Lovelace Institute (Ada Lovelace Institute 2022), and identified four key differences. First, our template enhances cross-company comparisons by aligning closely with legislation rather than relying on framework-specific biases or company-specific risks. Second, unlike existing templates which focus heavily on risks, ours equally emphasizes both risks and benefits. Third, there is variation in the extent of guidance provided for completing the templates. NIST’s (National Institute of Standards and Technology 2023b), Microsoft’s (Microsoft 2022a), and the Ada Lovelace Institute’s (Ada Lovelace Institute 2022) offer guidebooks on how to produce a report. Rather than leaving stakeholders with a blank report and a guide, we offer a guidebook through our 32 simple statements that alleviate potential anxiety associated with an empty page and facilitate the identification of risks and benefits. Finally, whereas other templates have separate sections for legal and compliance issues, ours integrates these aspects throughout.
@@
Improving stakeholder engagement. Our report template can be utilized by individual team members or as a group. For teams that are new to impact assessments, working through this framework can be an educational experience that helps in building understanding and skills related to responsible AI practices. Future work on improving stakeholder engagement should explore alternative interactive tools that effectively balance various methods of eliciting system information and envisioning system impacts, such as divergent and convergent thinking styles (Selbst 2021).
@@
Generalizability. Our results are based on a pool of study participants who were familiar with the EU AI Act, the NIST framework and ISO standard. Including a broader range of roles beyond managers, designers, and researchers may yield different results. Future studies should include participants with different levels of knowledge about AI regulations. They should also evaluate how well impact assessment reports explain the risks and benefits of AI to everyone and find ways to help people understand important AI regulations and laws.
Propagating biases in impact assessment. AIIAs, like other responsible AI tools, have inherent biases from their design choices (e.g., excluding potential users (Moss et al. 2021)). The quality of our report template depends on the accuracy and completeness of user-provided information. Using biased or incomplete data can lead to incorrect assessments, making problems seem smaller than they really are because people might be afraid to report negative impacts. To fix this, future research should include external perspectives by holding workshops, involving independent experts and marginalized groups, and regularly checking the results with new team members (Raji et al. 2020; Ada Lovelace Institute 2022).
Automated tools pre-populating the template. Automated tools can help make gathering data for templates easier and less prone to mistakes. Large Language Models (LLMs) can help fill out impact assessment reports by generating lists of AI users and subjects (Buçinca et al. 2023), identifying intended and unintended uses (Wang et al. 2024; Herdel et al. 2024), and listing potential risks and benefits (Constantinides et al. 2024b; De Miguel Velazquez et al. 2024). Bogucka et al. (2024) have recently proposed a semi-automatic system that collects input from stakeholders about an AI system’s use , uses LLMs to find additional risks, mitigation strategies, and benefits, and pre-fills reports for experts to review.
@@ Conclusion
In this study, we, the authors, are based in a Western country and engage predominantly in industry and academic research during the 21st century. Our team includes two women and two men from Southern, Eastern, and Central Europe, representing a variety of ethnic backgrounds. Our collective expertise spans several fields such as human-computer interaction, ubiquitous computing, software engineering, artificial intelligence, natural language processing, data visualization, and digital humanities. Our positionality may influence the inherent subjectivity in formulating our research, choosing our methods, structuring our co-design and evaluation sessions, interpreting and analyzing data, and addressing the needs of study participants in future iterations of the template (Frluckaj et al. 2022).
@@
The Artificial Intelligence Act (AI Act) is an upcoming regulatory framework for artificial intelligence in European Union. It analyses AI systems in various uses and categorizes them based on the risk they present to users into unacceptable risk, high-risk, limited risk and minimal risk uses.
@@
  • Article 9: Risk management system
    https://artificialintelligenceact.eu/article/9/
  • Article 10: Data and data governance
    https://artificialintelligenceact.eu/article/10/
  • Article 11: Technical Documentation and Annex IV: Technical documentation
    https://artificialintelligenceact.eu/article/11/ https://artificialintelligenceact.eu/annex/4/
  • Article 12: Record-keeping
    https://artificialintelligenceact.eu/article/12/
  • Article 13: Transparency and provision of information to deployers
    https://artificialintelligenceact.eu/article/13/
  • Article 14: Human oversight
    https://artificialintelligenceact.eu/article/14/
  • Article 15: Accuracy, robustness and cybersecurity
    https://artificialintelligenceact.eu/article/15/
  • Article 16: Obligations of providers of high-risk AI systems
    https://artificialintelligenceact.eu/article/16/
  • Article 18: Documentation keeping
    https://artificialintelligenceact.eu/article/18/
  • Article 27: Fundamental rights impact assessment for high-risk AI systems
    https://artificialintelligenceact.eu/article/27
  • Article 50: Transparency obligations for providers and users of certain AI systems
    https://artificialintelligenceact.eu/article/50/
  • Article 53: Obligations for providers of general-purpose AI models
    https://artificialintelligenceact.eu/article/53/
  • Article 55: Obligations for providers of general-purpose AI models with systemic risk
    https://artificialintelligenceact.eu/article/55/
  • Article 72: Post-market monitoring by providers and post-market monitoring plan for high-risk AI systems
    https://artificialintelligenceact.eu/article/72/
@@
The Artificial Intelligence Risk Management Framework is a voluntary guidance framework for organizations that design, develop, deploy, or use AI systems, aimed at helping them manage the risks associated with these systems. The framework is divided into two parts. The first part discusses how organizations can frame the risks related to AI and outlines the characteristics of trustworthy AI systems. These characteristics include being: valid and reliable, safe, secure and resilient, accountable and transparent, explainable and interpretable, privacy-enhanced, and fair with harmful bias managed.
The second part, the core of the framework, describes four specific functions to help organizations address the risks of AI systems in practice:
@@
  • Manage – prioritize risks and act upon their potential impact.
@@
This standard is directed to organizations that use, develop, monitor or provide products or services that utilize AI. Organizations shall establish a process to assess and document the potential consequences for individuals, groups of individuals, and societies that may result from the AI system throughout its life cycle.
Specifically, the organization should assess whether an AI system affects:
  • the legal position or life opportunities of individuals;
  • the physical or psychological well-being of individuals;
  • universal human rights;
  • societies.
Organization should consider documenting:
@@
V1 The primary concern with the first iteration with AI practitioners was the lack of evaluation details when the system is deployed in different contexts. P1 mentioned that “I didn’t see anything about how we were going to evaluate whether it was working in deployment […] some sort of details is needed about system’s use in actual deployment and what are the risks”. To resolve this issue, we added a new subsection “System evaluation”, that can be expanded with detailed information on the outcomes of evaluation activities at each stage of the system’s lifecycle: development, deployment, and use (Figure 5).
@@
First, to clarify the system’s use, we divided its description into five subsections matching the risk assessment components identified by Golpayegani (Goodell et al. 2021): purpose, capability, domain, user, and subject. Such a division provides all necessary information needed for conducting the risk assessment based on the EU AI Act. Users could specify these components using external dictionaries, such as the Vocabulary of AI Risks (Golpaygani, Pandit, and Lewis 2023), which contains lists of pre-defined descriptions for each of these concepts sourced directly from the Act. Second, after the ”Evaluation” subsection, we added a sub-section to list system’s use limitations. This subsection can cover for example the reasonably foreseeable misuse: the use of an AI system in a way that is not in accordance with its intended use, but which may result from reasonably foreseeable human behaviour or interaction with other systems, including other AI systems.
@@
V3 The primary concerns with the third iteration were a lack of information about the governance of the use and a lack of a systematic way of reporting risks, mitigations, and benefits of this use based on the origins of the risks or benefits and whom they can impact. For example, P10, wanted “to see a bit more about the actual system to understand the system’s transparency and its broader impact on society”. Additionally, E04 stated that when she “tried to explain to our teams that we need to surface risks for organization, individuals, groups of individuals, and societies, they found it too complicated and too remote to how currently we think about our tools and business.”.
@@
To resolve the first issue, we added to the template the final section with risk reporting methods (e.g., dedicated email, phone number, the registered office) and compliance certifications. To resolve the second issue, we proposed to divide the “Risks”, “Mitigations”, and “Benefits” sections into three subsections each, aligning them with a three-layered framework for evaluating sociotechnical harms (Weidinger et al. 2023) for risks stemming from the system’s capability (i.e., the technical components of the AI system), human interaction (i.e., experiences of people interacting with the AI system), and systemic impact (i.e., societal, economic, and environmental impacts of the AI system’s use).
@@
V4 The primary concern with the fourth iteration was the lack of scaffolding elements supporting populating the template. To provide such elements for each section of the template, we examined guidelines and questions proposed in documentation standards (Selbst 2021; Gebru et al. 2021; Holland et al. 2020; Bender and Friedman 2018; Mitchell et al. 2019; Sokol and Flach 2020; Raji et al. 2020), questionnaires grounded in regulations (The High-Level Expert Group on Artificial Intelligence 2020; Constantinides et al. 2024a), and checklists (Madaio et al. 2020; Golpayegani, Pandit, and Lewis 2023; Skoric 2023; National Institute of Standards and Technology 2023b) that structure ad-hoc practices in responsible AI. This resulted in a list of 32 statements that allow to systematically gather information about the system use, system components and data, the team involvement, and the risks, mitigations and benefits related to system’s use (Table 3).
@@
Figure 6: Impact assessment report for a meeting companion—an AI-based system aimed at monitoring employee behaviour during company meetings to improve meetings experience.
Figure 7: Impact assessment baseline for a meeting companion—an AI-based system aimed at monitoring employee behaviour during company meetings to improve meetings experience.
@@
Table 3 outlines 32 statements to systematically collect information for the impact assessment report from stakeholders. Each guideline is phrased in simple, actionable language for easy understanding by both technical and non-technical stakeholders and is accompanied by a practical example.
Statements 1–6 (System’s Use). To collect information about the system’s use, we considered five statements proposed by Golpayegani et al.  (Golpayegani, Pandit, and Lewis 2023) and one by Constantinides et al. (Constantinides et al. 2024a). These statements are about the system’s operational sector, users and subjects, thus providing a comprehensive description of the system’s scope of use and compliance to date. Together, these six statements are also useful for performing subsequent risk assessments of the use of the system  (Golpayegani, Pandit, and Lewis 2023). These statements are mapped to Section 1.1 of the final template (Figure 3).
@@
Statements 26 - 27 (Teams). These statements concern the teams behind the system use. We considered two statements from the Responsible AI guidelines (Constantinides et al. 2024a), which are about team diversity and training on ethical values and regulations. These statements are mapped to Section 1.4 of Figure 3.
@@
Statements 31 - 32 (Benefits). These statements are about the potential benefits of the system’s use for individuals, communities, organizations, and the planet. We sourced these statements from the NIST’s AI RMF (National Institute of Standards and Technology 2023a, b) and the ALTAI (The High-Level Expert Group on Artificial Intelligence 2020). They are mapped to Section 4 of the final template (Figure 3).
Table 3: List of 32 statements to systematically collect information for the impact assessment report from stakeholders.
ID Statement Source
SYSTEM’S USE
S1 Define the purpose of using the system (Golpayegani, Pandit, and Lewis 2023; Mitchell et al. 2019; Madaio et al. 2020; Raji et al. 2020)
Define purpose as job profile matching, assessing personality traits, managing road traffic
S2 Specify technical capability of the system (Golpayegani, Pandit, and Lewis 2023)
Specify the capability of image recognition, question answering, sentiment analysis
S3 Indicate the domain, industry or sector in which the system will operate (Golpayegani, Pandit, and Lewis 2023)
Indicate education, employment, public service
S4 Identify the individual subject affected by the use of the system (Golpayegani, Pandit, and Lewis 2023; Madaio et al. 2020; Raji et al. 2020)
List any individual, group or organization impacted by the system’s decisions
S5 Identify the entity or individual responsible for using and operating the system (Golpayegani, Pandit, and Lewis 2023; Mitchell et al. 2019; Raji et al. 2020)
List any organization, corporation, public authority, agency
S6 Obtain approval from an Ethics Committee or similar body for the system’s use (Constantinides et al. 2024a; Mitchell et al. 2019; Raji et al. 2020)
Confirm system compliance with sustainability goals
SYSTEM
Components (including models and monitoring techniques)
S7 Provide basic information about each model (Mitchell et al. 2019; Madaio et al. 2020)
Specify model date, version, type, license, documentation
S8 Describe the intended and out-of-scope tasks for each model (Mitchell et al. 2019; Madaio et al. 2020)
Describe the general or specific tasks for the model, provide model warnings and disclaimers
S9 Select relevant factors for evaluating each model’s performance such as groups, instruments, and deployment environments (Mitchell et al. 2019; Madaio et al. 2020; Raji et al. 2020)
Consider the visual presentation of age and gender
S10 Select relevant metrics for evaluating the real-world impacts of each model (Mitchell et al. 2019; Madaio et al. 2020; Raji et al. 2020)
Consider decision thresholds, uncertainty levels, and variability rates
S11 Report evaluation metrics for various groups (Constantinides et al. 2024a; Mitchell et al. 2019; Madaio et al. 2020)
Evaluate false positive/negative, AUC, and feature importance across protected attributes such as age, gender, and ethnicity
S12 Redocument each model information and contractual requirements at every system update (Constantinides et al. 2024a; Madaio et al. 2020)
Update the model information when re-training the system
S13 Document all system components, including all models, to enable reproducibility and scrutiny (Constantinides et al. 2024a; Mitchell et al. 2019; Madaio et al. 2020)
Create UML diagrams, and specify model types, versions, hardware architecture
S14 Provide mechanisms for interpretable outputs and auditing of all system components (Constantinides et al. 2024a; Sokol and Flach 2020)
Output feature importance and provide human-understandable explanations
S15 Document the security of all system components in consultation with experts (Constantinides et al. 2024a)
Guard against adversarial attacks and unauthorized access
S16 Review the code for reliability of the system (Constantinides et al. 2024a; Madaio et al. 2020)
Manage version control using software
S17 Develop feedback mechanisms to update the system (Constantinides et al. 2024a; Madaio et al. 2020)
Provide contact email, feedback form, and notification of new knowledge extracted
S18 Continuously monitor metrics and utilize guardrails or rollbacks to ensure the system’s output stays within a desired range (Constantinides et al. 2024a; Madaio et al. 2020)
Validate against concept drift and test with diverse testers
S19 Ensure human control over the system, particularly for designers, developers, and end-users (Constantinides et al. 2024a; Madaio et al. 2020)
Include human in the loop with the ability to inspect data, models, and training methods
S20 Ensure safe system decommissioning (Constantinides et al. 2024a; Madaio et al. 2020)
Ensure decommissioned data is either deleted or restricted to authorized personnel
S21 Provide an environmental assessment of the system (Constantinides et al. 2024a)
Report the number of GPU hours used in training and deployment
Data
S22 Provide information about the training data, including any measurement errors in input data and their associated assumptions (Constantinides et al. 2024a; Holland et al. 2020; Mitchell et al. 2019; Gebru et al. 2021)
Describe the distribution of the data over key factors
S23 Compare the quality, representativeness, and fit of training and testing datasets with the system uses (Constantinides et al. 2024a; Holland et al. 2020; Madaio et al. 2020; Gebru et al. 2021)
Report dataset details such as demographics, and data provenance
S24 Protect sensitive variables in training/testing datasets (Constantinides et al. 2024a)
Protect sensitive data and use techniques such as k-anonymity and differential privacy
S25 Ensure compliance with agreements and legal requirements when handling data (Constantinides et al. 2024a; Gebru et al. 2021)
Create data sharing and non-disclosure agreements and secure servers
TEAMS
S26 Ensure team diversity (Constantinides et al. 2024a; National Institute of Standards and Technology 2023b)
Consider diversity in gender, neurotypes, personality traits, and thinking styles
S27 Train team members on ethical values and regulations (Constantinides et al. 2024a)
Train on privacy regulations, ethical issues, and raising concerns
RISKS AND MITIGATIONS
S28 Identify potential harms and risks associated with system use (Constantinides et al. 2024a; Mitchell et al. 2019; Madaio et al. 2020; Raji et al. 2020)
Prevent privacy violation, discrimination, and adversarial attacks
S29 Provide mechanism(s) for incentivizing reporting of system harms and risks (Constantinides et al. 2024a; Madaio et al. 2020)
Provide contact emails and feedback form to raise concerns
S30 Develop strategies to mitigate identified harms or risks for each system use (Constantinides et al. 2024a; Mitchell et al. 2019; Madaio et al. 2020; Raji et al. 2020)
Use stratified sampling and safeguards against adversarial attacks during training
BENEFITS
S31 Identify the benefits of the system’s use for individuals, communities, organizations, and society (National Institute of Standards and Technology 2023b)
Describe how the system benefits individuals from marginalized communities (The High-Level Expert Group on Artificial Intelligence 2020)
S32 Identify the benefits of the system’s use for the planet (National Institute of Standards and Technology 2023b)
Describe how the system provides benefits to natural and cultural environments (The High-Level Expert Group on Artificial Intelligence 2020)
@@
  • How old are you?
  • What is your gender? [Male, Female, Non-binary, Prefer not to say, Open-ended option]
  • How many years of experience do you have in AI systems?
@@
  • What is domain or sector of your work? (e.g., health, energy, education, finance, technology, food)
  • What is your current role?
  • What kinds of AI systems do you work on? (e.g., machine learning, computer vision, NLP, game theory, robotics)

### v2

@@ Discussion
Differences from Existing Templates
We compared our template with the Algorithmic Impact Assessment template from NIST (National Institute of Standards and Technology 2023b), Microsoft’s Responsible AI Impact Assessment Template (Microsoft 2022b), Credo AI’s Standardized Risk Profile (Sherman and Eisenberg 2024), and the Algorithmic Impact Assessment from the Ada Lovelace Institute (Ada Lovelace Institute 2022), and identified four key differences. First, our template enhances cross-company comparisons by aligning closely with legislation rather than relying on framework-specific biases or company-specific risks. Second, unlike existing templates which focus heavily on risks, ours equally emphasizes both risks and benefits. Third, there is variation in the extent of guidance provided for completing the templates. NIST’s (National Institute of Standards and Technology 2023b), Microsoft’s (Microsoft 2022a), and the Ada Lovelace Institute’s (Ada Lovelace Institute 2022) offer guidebooks on how to produce a report. Rather than leaving stakeholders with a blank report and a guide, we offer a guidebook through our 32 simple statements that alleviate potential anxiety associated with an empty page and facilitate the identification of risks and benefits (Bogucka et al. 2024b). Finally, whereas other templates have separate sections for legal and compliance issues, ours integrates these aspects throughout.
@@
Improving stakeholder engagement. Our report template can be utilized by individual team members or as a group. For teams that are new to impact assessments, working through this template can be an educational experience that helps in building understanding and skills related to responsible AI practices. Future work on improving stakeholder engagement should explore alternative interactive tools that effectively balance various methods of eliciting system information and envisioning system impacts, such as divergent and convergent thinking styles (Selbst 2021).
@@
Generalizability. Our results are based on a pool of study participants who were familiar with the EU AI Act, the NIST framework and ISO standards. Including a broader range of roles beyond managers, designers, and researchers may yield different results. Future studies should include participants with different levels of knowledge about AI regulations. They should also evaluate how well impact assessment reports explain the risks and benefits of AI to everyone and find ways to help people understand important AI regulations and laws.
Propagating biases in impact assessment. AIIAs, like other responsible AI tools, have inherent biases from their design choices, such as excluding potential users (Moss et al. 2021). The quality of our report template depends on the accuracy and completeness of user-provided information. Using biased or incomplete data can lead to incorrect assessments, making problems seem smaller than they really are because people might be afraid to report negative impacts. To fix this, future research should include external perspectives by holding workshops, involving independent experts and marginalized groups, and regularly checking the results with new team members (Raji et al. 2020; Ada Lovelace Institute 2022).
Automated tools pre-populating the template. Automated tools can help make gathering data for templates easier and less prone to mistakes. Large Language Models (LLMs) can help fill out impact assessment reports by generating lists of AI users and subjects (Buçinca et al. 2023), identifying intended and unintended uses (Wang et al. 2024; Herdel et al. 2024), and listing potential risks and benefits (Constantinides et al. 2024b; De Miguel Velazquez et al. 2024). Bogucka et al. (2024a) have recently proposed a semi-automatic system that collects input from stakeholders about an AI system’s use , uses LLMs to find additional risks, mitigation strategies, and benefits, and pre-fills reports for experts to review.
@@ Conclusion
In this study, we, the authors, are based in United Kingdom and engage predominantly in industry and academic research during the 21st century. Our team includes two women and two men from Southern, Eastern, and Central Europe, representing a variety of ethnic backgrounds. Our collective expertise spans several fields such as human-computer interaction, ubiquitous computing, software engineering, artificial intelligence, natural language processing, data visualization, and digital humanities. Our positionality may influence the inherent subjectivity in formulating our research, choosing our methods, structuring our co-design and evaluation sessions, interpreting and analyzing data, and addressing the needs of study participants in future iterations of the template (Frluckaj et al. 2022).
@@
The Artificial Intelligence Act (AI Act) is a regulatory framework for artificial intelligence in European Union. It analyses AI systems in various uses and categorizes them based on the risk they present to users into unacceptable risk, high-risk, limited risk and minimal risk uses.
@@
  • Art. 9: Risk management system
  • Art. 10: Data and data governance
  • Art. 11: Technical Documentation and
  • Art. 12: Record-keeping
  • Art. 13: Transparency and provision of information to deployers
  • Art. 14: Human oversight
  • Art. 15: Accuracy, robustness and cybersecurity
  • Art. 17: Quality management system
  • Art. 16: Obligations of providers of high-risk AI systems
  • Art. 18: Documentation keeping
  • Art. 27: Fundamental rights impact assessment for high-risk AI systems
  • Art. 50: Transparency obligations for providers and users of certain AI systems
  • Art. 53: Obligations for providers of general-purpose AI models
  • Art. 55: Obligations for providers of general-purpose AI models with systemic risk
  • Art. 72: Post-market monitoring by providers and post-market monitoring plan for high-risk AI systems
Source: https://artificialintelligenceact.eu
@@
The Artificial Intelligence Risk Management Framework (National Institute of Standards and Technology 2023a) is a voluntary guidance framework for organizations that design, develop, deploy, or use AI systems, aimed at helping them manage the risks associated with these systems. The framework is divided into two parts. The first part discusses how organizations can frame the risks related to AI and outlines the characteristics of trustworthy AI systems. These characteristics include being: valid and reliable, safe, secure and resilient, accountable and transparent, explainable and interpretable, privacy-enhanced, and fair. The second part, the core of the framework, describes four specific functions to help organizations address the risks of AI systems in practice:
@@
  • Manage – prioritize risks and act upon their impact.
@@
This standard applies to organizations using, developing, monitoring, or providing AI products or services. Organizations shall establish a process to assess and document the potential consequences for individuals, groups of individuals, and societies that may result from the AI system through its life cycle. Specifically, the organization should assess whether an AI system affects the legal position or life opportunities of individuals, the physical or psychological well-being of individuals, universal human rights, and societies, and document:
@@
V1 The primary concern with the first iteration with AI practitioners was the lack of evaluation details when the system is deployed in different contexts. P1 mentioned that ‘more details are needed about system’s use in actual deployment and what are the risks”. To resolve this issue, we added a new subsection “System evaluation”, which details evaluation outcomes at each stage of the system’s lifecycle: development, deployment, and use (Figure 4).
@@
First, to clarify the system’s use, we divided its description into five subsections matching the risk assessment components identified by  Golpayegani, Pandit, and Lewis (2023): purpose, capability, domain, user, and subject. Such a division provides all necessary information needed for conducting the risk assessment based on the EU AI Act. Users could specify these components using external dictionaries, such as the Vocabulary of AI Risks (Golpaygani, Pandit, and Lewis 2023), which contains lists of descriptions for each of these concepts sourced directly from the Act.
Second, after the ”Evaluation” subsection, we added a sub-section to list system’s use limitations. This subsection can cover for example the reasonably foreseeable misuse: the use of an AI system in a way that is not in accordance with its intended use, but which may result from reasonably foreseeable human behaviour or interaction with other systems, including other AI systems.
@@
V3 The primary concerns with the third iteration were a lack of information about the governance of the use and a lack of a systematic way of reporting risks, mitigations, and benefits of this use based on the origins of the risks or benefits and whom they can impact. For example, P10, wanted “to see a bit more about the actual system to understand the system’s transparency and its broader impact on society”. Additionally, E04 stated that when she “tried to explain to our teams that we need to surface risks for organization, individuals, and groups of individuals, they found it too complicated and too remote to how currently we think about our business.”.
@@
To resolve the first issue, we added to the template the final section with risk reporting methods (e.g., dedicated email, the registered office) and compliance certifications. To resolve the second issue, we proposed to divide the “Risks”, “Mitigations”, and “Benefits” sections into three subsections each, aligning them with a three-layered framework for evaluating sociotechnical harms (Weidinger et al. 2023) for risks stemming from the system’s capability (i.e., the technical components of the AI system), human interaction (i.e., experiences of people interacting with the AI system), and systemic impact (i.e., societal, economic, and environmental impacts of the AI system’s use).
@@
V4 The primary concern with the fourth iteration was the lack of scaffolding elements supporting populating the template. To provide such elements for each section of the template, we examined guidelines and questions proposed in documentation standards (Selbst 2021; Gebru et al. 2021; Holland et al. 2020; Bender and Friedman 2018; Mitchell et al. 2019; Sokol and Flach 2020; Raji et al. 2020), questionnaires grounded in regulations (The High-Level Expert Group on Artificial Intelligence 2020; Constantinides et al. 2024a), and checklists (Madaio et al. 2020; Golpayegani, Pandit, and Lewis 2023; Skoric 2023; National Institute of Standards and Technology 2023b) that structure ad-hoc practices in responsible AI. This resulted in a list of 32 statements that allow to systematically gather information about the system use, system components and data, the team involvement, and the risks, mitigations and benefits related to system’s use (Bogucka et al. 2024b).
@@
Our project page (Bogucka et al. 2024b) outlines 32 statements to systematically collect information for the impact assessment report from stakeholders. Each guideline is phrased in simple, actionable language for easy understanding by both technical and non-technical stakeholders and is accompanied by a practical example.
Statements 1–6 (System’s Use). To collect information about the system’s use, we considered five statements proposed by Golpayegani, Pandit, and Lewis (2023) and one by Constantinides et al. (2024a). These statements are about the system’s operational sector, users and subjects, thus providing a comprehensive description of the system’s scope of use and compliance to date. Together, they are also useful for subsequent risk assessments of the use of the system  (Golpayegani, Pandit, and Lewis 2023). These statements are mapped to Section 1.1 of the final template (Figure 2).
@@
Statements 26 - 27 (Teams). These statements concern the teams behind the system use. We considered two statements from the Responsible AI guidelines (Constantinides et al. 2024a), which are about team diversity and training on ethical values and regulations. They are mapped to Section 1.4 of Figure 2.
@@
Statements 31 - 32 (Benefits). These statements address the benefits of the system’s use for individuals, communities, organizations, and the planet. We sourced these statements from the NIST’s AI RMF (National Institute of Standards and Technology 2023a, b) and the ALTAI (The High-Level Expert Group on Artificial Intelligence 2020). They are mapped to Section 4 of the final template (Figure 2).
@@
  • How old are you? What is your gender?
  • How many years of experience do you have in AI?
@@
  • What is domain or sector of your work? (e.g., health)
  • What is your current role? What kinds of AI systems do you work on (e.g., machine learning, computer vision)?
Figure 5: Impact assessment report for a meeting companion—an AI-based system aimed at monitoring employee behaviour during company meetings to improve meetings experience (Bogucka et al. 2024b).
Figure 6: Impact assessment baseline for a meeting companion—an AI-based system aimed at monitoring employee behaviour during company meetings to improve meetings experience (Bogucka et al. 2024b).